Donald Martin is planning to invest $29,000 today in a mutual fund that will provide a return of 10 percent each year. What will be the value of the investment in 10 years? (If you solve this problem with algebra round intermediate calculations to 6 decimal places, in all cases round your final answer to the nearest penny.) Value of investment after 10 years

Debra King received a graduation present of $1,000 that she is planning on investing in a mutual fund that earns 15.50 percent each year. How much money will she have in three years? (If you solve this problem with algebra round intermediate calculations to 6 decimal places, in all cases round your final answer to the nearest penny.)

Value of investment after 3 years
Answer #1

Value of investment after 10 years is computed as shown below:

Future value = Present value ( 1 +r )n

= $ 29,000 (1 + 0.10 )10

= $ 75,218.53 Approximately

Money that she will have in three years is computed as shown below:

Future value = Present value ( 1 +r )n

= $ 1,000 ( 1 + 0.1550 )3

= $ 1,540.80 Approximately

Feel free to ask in case of any query relating to this question
Answer #2

SOLUTION :


Value of Donald Martin’s investment after10 years 

= A(1+r)^n

= 29000*(1 + 0.10)^10

= 75218.53 ($) (ANSWER)


Value of Debra King’s investment in 3 years 

= A(1+r)^n

= 1000*(1 + 0.155)^3

= 1540.80 ($) (ANSWER)

answered by: Tulsiram Garg